Find the surface area of the following rectangular prism
L=12, H=5, W= 2

Answer by Mathtut(3670)

there are 6 areas to add together.....opposite sides have equal areas
:
so find the areas of the three different sides and multiply each by 2 and then add them all together
:
2(2*5)+2(2*12)+2(5*12)=20+24+120=164 sq units